# Linear Bandits with Partially Observable Features

### v2

@@ 4.1 Feature Vector Augmentation with Orthogonal Projection
In the linear bandit framework, accurate estimation of the unknown reward parameter $\boldsymbol{\theta}_{\star}$ contributes to optimal decision making. However, in our problem setting, the learner lacks access to the latent reward component—namely, the last term in Eq. 3. Consequently, without any specific treatment to compensate for the absence of $\boldsymbol{\theta}_{\star}^{(u)}$, the agent fails to accurately estimate $\boldsymbol{\theta}_{\star}$, resulting in ineffective learning, as demonstrated in in Theorem 1.
That said, minimizing regret—that is, selecting the optimal arm—does not require recovering the true reward parameter $\boldsymbol{\theta}_{\star}$ entirely; rather, it suffices to estimate the $K$ expected rewards $\{\mathbf{z}_{a}^{\top}\boldsymbol{\theta}_{\star}:a\in[K]\}$. A straightforward approach is to ignore the features altogether and apply MAB algorithms such as UCB1 (Auer et al., 2002), which are known to achieve a regret bound of $\widetilde{O}(\sqrt{KT})$. However, these algorithms tend to incur higher regret than those leveraging features, particularly when the number of arms is significantly larger than the dimension of the feature vectors, i.e., $K\gg d_{z}$.
To tackle this dilemma, we propose a unified approach to handle all cases of partially observable features and efficiently estimate the expected rewards. Let $\mathbf{X}:=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{d\times K}$ represent a matrix that concatenates the observed part of the true features, and $\mathbf{U}:=(\mathbf{u}_{1}^{(u)},\ldots,\mathbf{u}_{K}^{(u)})\in\mathbb{R}^{d_{u}\times K}$ represent the matrix that concatenates the latent complements of the true features for each arm. We assume a set of $K$ vectors $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$ spans $\mathbb{R}^{d}$, without loss of generality. [^3] When $d>K$, we can apply singular value decomposition on $\mathbf{X}$ to reduce the feature dimension to $\bar{d}\leq K$ with $\mathrm{R}(\mathbf{X})=\bar{d}$. We define $\mathbf{P}_{\mathbf{X}}:=\mathbf{X}^{\top}(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}$ as the projection matrix onto the row space of $\mathbf{X}$, denoted by $\mathrm{R}(\mathbf{X})$. Then the vector of rewards for all arms, $\mathbf{Y}_{t}=(y_{1,t},\ldots,y_{K,t})^{\top}$, is now decomposed as:
@@
where the first term corresponds to the reward projected onto $\mathrm{R}(\mathbf{X})$; while the second term is the projected rewards onto $\mathrm{R}(\mathbf{X})^{\perp}$, the subspace of $\mathbb{R}^{K}$ perpendicular to $\mathrm{R}(\mathbf{X})$. We write the projected parameter as ${\boldsymbol{\mu}_{\star}^{(o)}}:=\boldsymbol{\theta}_{\star}^{(o)}+(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}$.
To handle the second term in Eq. 4, we consider a set of row vector basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}\in\mathrm{R}(\mathbf{X})^{\perp}$, where $\mathbf{b}_{i}\in\mathbb{R}^{K}$ for $i\in[K-d]$. Given the set, there exist coefficients ${\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}\in\mathbb{R}$ to express the reward projection as:
@@
We denote the number of nonzero coefficients:
@@
Note that $d_{h}=0$ for any basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$ when the latent feature space is completely included in the observed feature space, i.e., $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$. In this case, $(\mathbf{I}_{K}-\mathbf{P}_{X})\mathbf{U}^{\top}=\mathbf{0}_{K\times d_{u}}$, which means that the projected rewards onto $\mathrm{R}(\mathbf{X})^{\perp}$ can be linearly expressed by the observed features. If $\mathrm{R}(\mathbf{U})\supseteq\mathrm{R}(\mathbf{X})$, on the other hand, then $d_{h}=K-d$ for any basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$.
@@
In the other cases, the quantity $d_{h}$ depends on the choice of the basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$, and tends to be smaller when a larger portion of $\mathrm{R}(\mathbf{U})$ can be expressed within $\mathrm{R}(\mathbf{X})$. For any choice of the basis, our algorithm achieves $\widetilde{O}(\sqrt{(d+d_{h})T})$ regret without prior knowledge of $d_{h}$, which does not exceed $\widetilde{O}(\sqrt{KT})$ regret bound achieved by MAB algorithms ignoring features. Further details are provided in Section 5.2.
Let us define $\boldsymbol{\mu}_{\star}$ as $[({\boldsymbol{\mu}_{\star}^{(o)}})^{\top},({\boldsymbol{\mu}_{\star}^{(u)}})^{\top}]^{\top}\in\mathbb{R}^{K}$, where $\boldsymbol{\mu}_{\star}^{(u)}=[{\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}]^{\top}$, then the reward for each $a\in[K]$ can be modified as follows:
$\begin{split}y_{a,t}&=\mathbf{e}_{a}^{\top}\mathbf{Y}\\ &=\mathbf{e}_{a}^{\top}[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\\ &=[{\mathbf{x}_{a}}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}.\end{split}$ (7)
@@
The decomposition in Eq. 7 implies that Eq. 4 takes the form $\mathbf{Y}_{t}=[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\boldsymbol{1}_{K}$. Note that $\mathbf{e}_{a}\in\mathbb{R}^{K}$ is a Euclidean basis, with elements all zero except for $1$ in the $a$-th coordinate. With this modification, the rewards are now represented as a linear function of the augmented feature vectors: $\widetilde{\mathbf{x}}_{a}:=[\mathbf{x}_{a}^{\top}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]^{\top}\in\mathbb{R}^{K}$, without any misspecification error. A toy example illustrating our strategy is shown in Figure 1.
@@
Figure 1: Illustration comparing OLB algorithms and our approach in estimating rewards of $K=3$ arms. OLB algorithms find estimates within $\mathrm{R}(\mathbf{X})$ thus accumulating errors from unobserved features. However, our approach utilizes projection of the latent reward, $\mathbf{b}^{\top}\widehat{\boldsymbol{\mu}}_{t}^{(u)}$, onto the orthogonal complement of $\mathrm{R}(\mathbf{X})$, enabling reward estimation in $\mathbb{R}^{K}$. Note that $\widehat{\boldsymbol{\mu}}_{t}$ is the estimator of the parameter for observed features.
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathrm{R}(\mathbf{X})$$\mathbf{q}_{2}$$\mathbf{X}^{\top}\widehat{\boldsymbol{\mu}}_{t}$
@@
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathbf{q}_{2}$($\mathrm{R}(\mathbf{X})$$\mathbf{b}^{\top}\widehat{\boldsymbol{\mu}}_{t}^{(u)}$$\mathbf{X}^{\top}\widehat{\boldsymbol{\mu}}_{t}$
@@
The dimension of the augmented feature vectors $\{\widetilde{\mathbf{x}}_{a}:a\in[K]\}$ is $K\geq d$. While applying SupLinUCB (Chu et al., 2011) yields $\widetilde{O}(\sqrt{dT})$ regret bound, it is known to lack practicality as it computes $\log T$ distinct batches and estimators requiring the knowledge of $T$ and, more critically, discards a significant portion of samples in each parameter update. We present an efficient algorithm that employs the doubly robust ridge estimator and achieves $\widetilde{O}(\sqrt{KT})$ regret bound (see Appendix C). However, when $K>d$ and $d_{u}=0$, the regret is high compared to the linear bandits with conventional features. To address this challenge, we propose a novel estimation strategy in the following section that eliminates the dependence on $K$ in the regret bound.
@@ 4.2 Doubly Robust Lasso Estimator
In Eq. 7, the parameter $\boldsymbol{\mu}_{\star}$ is sparse, with its sparsity depending on the dimension of the subspace that captures the inaccessible portion of the reward—$d_{h}$ defined in Eq. 6. Recall that $\boldsymbol{\mu}_{\star}^{(u)}$ contains the coefficients used to express the projection of the reward, as represented in Eq. 5, and that only $d_{h}$ basis vectors are required to express this projection. Therefore, $\boldsymbol{\mu}_{\star}^{(u)}$ contains at most $d_{h}$ nonzero entries.
@@
$\begin{split}\check{\boldsymbol{\mu}}^{L}_{t}&:=\mathop{\mathrm{argmin}}_{\boldsymbol{\mu}}\sum_{\tau=1}^{t}(y_{a_{\tau},\tau}-\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}\boldsymbol{\mu})^{2}\\ &\qquad+2\widetilde{\sigma}_{\max}\sigma\sqrt{2pt\log\frac{2Kt^{2}}{\delta}}\|\boldsymbol{\mu}\|_{1},\end{split}$ (8)
@@
where $p$ is the coupling probability used to define the multinomial distribution for pseudo-action sampling (as defined in Eq. 9), and $\delta$ is the confidence parameter of the algorithm. Note that $\widetilde{\sigma}^{2}_{\max}:=\max_{a\in[K]}\mathbf{e}_{a}^{\top}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\mathbf{e}_{a}$, i.e., the largest diagonal entry of the Gram matrix constructed from the augmented feature vectors $\widetilde{\mathbf{x}}_{a}$ over $\mathcal{A}$.
@@
For the estimator in Eq. 8 to correctly identify the zero entries in $\boldsymbol{\mu}_{\star}$, the compatibility condition is required to be satisfied (van de Geer & Bühlmann, 2009). Although the compatibility condition does not, in general, require a positive minimum eigenvalue, in our setting the Gram matrix $\lambda_{\min}(t^{-1}\sum_{s=1}^{t}\widetilde{x}_{a_{s}}\widetilde{x}_{a_{s}}^{\top})$ has a strictly positive minimum eigenvalue. Therefore, the compatibility condition is implied without any additional assumption. However, ensuring a sufficiently large minimum eigenvalue typically requires collecting a large number of exploration samples, which in turn increases regret. Achieving this with fewer exploration samples remains a key challenge in bandit literature, as the minimum eigenvalue affects the convergence rate of the estimator and, consequently, the regret bound (Soare et al., 2014; Kim et al., 2021).
@@
We introduce a doubly robust (DR) estimator that employs the Gram matrix constructed from the augmented feature vectors over the entire action space, $\sum_{s=1}^{t}\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$, instead of the Gram matrix based only on the vectors chosen at each round, i.e., $\sum_{s=1}^{t}\widetilde{\mathbf{x}}_{a_{s}}\widetilde{\mathbf{x}}_{a_{s}}^{\top}$. The DR estimation originates from the statistical literature on missing data (Bang & Robins, 2005), where “doubly robust” refers to estimators that remain consistent if either the imputation model or the observation probability model is correctly specified (Kim et al., 2021). In the bandit setting, at each decision round $t\in[T]$, only the reward corresponding to the selected arm is observed, while the $K-1$ unselected rewards are missing. Hence, DR estimation imputes these $K-1$ missing rewards and incorporates the associated feature vectors into the estimation process. Since the observation probabilities are determined by the policy—which is known to the learner—the DR estimator remains robust to errors in the reward estimation. While Kim & Paik (2019) proposed a DR Lasso estimator under IID features satisfying the compatibility condition, we propose a novel DR Lasso estimator that does not rely on such assumptions.
@@
We improve the DR estimation by incorporating resampling and coupling methods. For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote the set of exploration rounds such that for any $\tau\in\mathcal{E}_{t}$, the action $a_{\tau}$ is sampled uniformly over $[K]$. The set $\mathcal{E}_{t}$ is constructed recursively. Starting with $\mathcal{E}_{0}=\emptyset$, we define
@@
$\mathcal{E}_{t}=\begin{cases}\mathcal{E}_{t-1}\cup\{t\}&\mathrm{if}\quad|\mathcal{E}_{t}|\leq C_{e}\log\dfrac{2Kt^{2}}{\delta},\\ \mathcal{E}_{t-1}&\mathrm{otherwise}.\end{cases}\vspace{-0.2cm}$
@@
Here, $C_{e}$ is defined as $(8K)^{3}\widetilde{\sigma}_{\min}^{-2}\widetilde{\sigma}_{\max}^{2}(1-p)^{-2}$, where $\widetilde{\sigma}_{\min}^{2}:=\lambda_{\min}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})$ denotes the minimum eigenvalue of the augmented Gram matrix over $\mathcal{A}$.
@@
At each round $t$, when $t\notin\mathcal{E}_{t}$, the algorithm selects an action $a_{t}$ according to an $\epsilon_{t}$-greedy policy. Then, a pseudo-action $\widetilde{a}_{t}$ is generated from a multinomial distribution:
@@
When the DR estimation is applied under the $\epsilon_{t}$-greedy policy, the pseudo-reward (Eq. 10) involves inverse probability term $\epsilon_{t}^{-1}:=\sqrt{t}$, which can cause its variance to grow unbounded over time. To mitigate this issue, we propose a resampling and coupling strategy: by coupling the $\epsilon_{t}$-greedy policy with the multinomial distribution Eq. 9, we ensure that each inverse probability weight $\phi_{a,t}^{-1}$, for $a\in[K]$, remains bounded by $O(K)$. This coupling yields a lower bound on the observation probability, thereby reducing the variance of the DR pseudo-rewards (Eq. 10).
Moreover, one can show that the resampling succeeds with high probability for each round. Let $\mathcal{M}_{t}$ denote an event that the pseudo-action $\widetilde{a}_{t}$ matches the chosen action $a_{t}$ within a specified number of resamples. For a given $\delta^{\prime}\in(0,1)$, we set the maximum number of resamples to $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/p)$, then $\mathcal{M}_{t}$ occurs with probability at least $1-\delta^{\prime}/(t+1)^{2}$. This resampling allows the algorithm to further explore the action space in order to balance regret minimization with accurate reward estimation.
With the pseudo-actions coupled with the actual actions, we construct unbiased pseudo-rewards for all $a\in[K]$ as follows:

### v3

@@ 4.1 Feature Vector Augmentation with Orthogonal Projection
In the linear bandit framework, accurate estimation of the unknown reward parameter $\boldsymbol{\theta}_{\star}$ contributes to optimal decision-making. However, in our problem setting, the learner lacks access to the latent reward component—namely, the last term in Eq.˜3. Consequently, without an appropriate mechanism to compensate for the absence of $\boldsymbol{\theta}_{\star}^{(u)}$, the agent fails to accurately estimate $\boldsymbol{\theta}_{\star}$, resulting in ineffective learning, as demonstrated in Theorem˜1.
That said, minimizing regret—that is, selecting the optimal arm—does not require recovery of the true reward parameter $\boldsymbol{\theta}_{\star}$ entirely; rather, it suffices to estimate the $K$ expected rewards $\{\mathbf{z}_{a}^{\top}\boldsymbol{\theta}_{\star}:a\in[K]\}$. A straightforward approach is to ignore the features altogether and apply MAB algorithms such as UCB1 (Auer et al., 2002), which are known to achieve a regret bound of $\widetilde{O}(\sqrt{KT})$. However, these algorithms tend to suffer higher regret than feature-based algorithms, particularly when the number of arms is significantly larger than the dimension of the feature vectors, i.e., $K\gg d_{z}$.
To tackle this dilemma, we propose a unified approach that handles partial observability and ensures efficient estimation. Let us define $\mathbf{X}:=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{d\times K}$ as the matrix formed by concatenating the observed part of the true features, and $\mathbf{U}:=(\mathbf{u}_{1}^{(u)},\ldots,\mathbf{u}_{K}^{(u)})\in\mathbb{R}^{d_{u}\times K}$ as the matrix formed by concatenating their latent complements for each arm. Without loss of generality, we assume a set of $K$ vectors $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$ spans $\mathbb{R}^{d}$. [^3] When $d>K$, we can apply singular value decomposition (SVD) on $\mathbf{X}$ to reduce the feature dimension to $\bar{d}\leq K$ with $\mathrm{R}(\mathbf{X})=\bar{d}$. We denote by $\mathbf{P}_{\mathbf{X}}:=\mathbf{X}^{\top}(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}$ the projection matrix onto the row space of $\mathbf{X}$, denoted by $\mathrm{R}(\mathbf{X})$. Then the vector of rewards for all arms, $\mathbf{Y}_{t}=(y_{1,t},\ldots,y_{K,t})^{\top}$, is decomposed as:
@@
where the first term corresponds to the reward projected onto $\mathrm{R}(\mathbf{X})$, whereas the second term is the reward projected onto $\mathrm{R}(\mathbf{X})^{\perp}$, the subspace of $\mathbb{R}^{K}$ orthogonal to $\mathrm{R}(\mathbf{X})$. We reparametrize $\boldsymbol{\theta}_{\star}^{(o)}+(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}$ as $\boldsymbol{\mu}_{\star}^{(o)}$}, representing the parameter associated with the observed features.
To handle the second term in Eq.˜4, we consider a set of row vector bases $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}\in\mathrm{R}(\mathbf{X})^{\perp}$, where $\mathbf{b}_{i}\in\mathbb{R}^{K}$ for $i\in[K-d]$. Given the set, there exist coefficients ${\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}\in\mathbb{R}$ such that the term can be expressed as a linear combination:
@@
We define the number of nonzero coefficients as:
@@
Note that $d_{h}=0$ for any basis set $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$ when the latent feature space is completely included in the observed feature space, i.e., $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$. In this case, $(\mathbf{I}_{K}-\mathbf{P}_{\mathbf{X}})\mathbf{U}^{\top}=\mathbf{0}_{K\times d_{u}}$, which means that the projected rewards onto $\mathrm{R}(\mathbf{X})^{\perp}$ can be linearly expressed by the observed features. If $\mathrm{R}(\mathbf{U})\supseteq\mathrm{R}(\mathbf{X})$, on the other hand, then $d_{h}=K-d$ for any basis set $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$.
@@
In other cases, the quantity $d_{h}$ depends on the choice of the basis $\{\mathbf{b}_{1}^{\top},\ldots,\mathbf{b}_{K-d}^{\top}\}$, and tends to be smaller when a larger portion of $\mathrm{R}(\mathbf{U})$ is included within $\mathrm{R}(\mathbf{X})$. For any choice of the basis, our algorithm achieves $\widetilde{O}(\sqrt{(d+d_{h})T})$ regret without prior knowledge of $d_{h}$, which does not exceed the $\widetilde{O}(\sqrt{KT})$ regret bound achieved by MAB algorithms ignoring features. Further details are provided in Section˜5.2.
Similar to the reparametrization of $\boldsymbol{\mu}_{\star}^{(o)}$, we denote $\boldsymbol{\mu}_{\star}^{(u)}:=[\mu_{\star,1}^{(u)},\ldots,\mu_{\star,K-d}^{(u)}]^{\top}$. By concatenating $\boldsymbol{\mu}_{\star}^{(o)}$ and $\boldsymbol{\mu}_{\star}^{(u)}$, we define $\boldsymbol{\mu}_{\star}:=[({\boldsymbol{\mu}_{\star}^{(o)}})^{\top},({\boldsymbol{\mu}_{\star}^{(u)}})^{\top}]^{\top}\in\mathbb{R}^{K}$, so that the reward for each $a\in[K]$ can be rewritten as follows:
$y_{a,t}$ $=\mathbf{e}_{a}^{\top}\mathbf{Y}$
$=\mathbf{e}_{a}^{\top}[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}$
$=[{\mathbf{x}_{a}}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}.$ (7)
@@
The decomposition in Eq.˜7 implies that Eq.˜4 takes the form $\mathbf{Y}_{t}=[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\boldsymbol{1}_{K}$. Note that $\mathbf{e}_{a}\in\mathbb{R}^{K}$ denotes the $a$-th standard basis vector. With this modification, the rewards are now represented as a linear function of the augmented feature vectors: $\widetilde{\mathbf{x}}_{a}:=[\mathbf{x}_{a}^{\top}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]^{\top}\in\mathbb{R}^{K}$, without any misspecification error. A toy example illustrating our strategy is shown in Figure˜1.
@@
Figure 1: Illustration comparing OLB algorithms and our approach in estimating rewards of $K=3$ arms. OLB algorithms find estimates within $\mathrm{R}(\mathbf{X})$ thus accumulating errors from unobserved features. However, our approach utilizes the projection of the latent reward, $\mathbf{b}^{\top}\widehat{\boldsymbol{\mu}}_{t}^{(u)}$, onto the orthogonal complement of $\mathrm{R}(\mathbf{X})$, thereby enabling reward estimation in $\mathbb{R}^{K}$.
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathrm{R}(\mathbf{X})$$\mathbf{q}_{2}$$\mathbf{X}^{\top}\widehat{\boldsymbol{\mu}}_{t}^{(o)}$
@@
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathbf{q}_{2}$$\mathrm{R}(\mathbf{X})$$\mathbf{b}^{\top}\widehat{\boldsymbol{\mu}}_{t}^{(u)}$$\mathbf{X}^{\top}\widehat{\boldsymbol{\mu}}_{t}^{(o)}$
@@
In terms of regret minimization, while SupLinUCB (Chu et al., 2011) achieves a regret bound of $\widetilde{O}(\sqrt{dT})$, it is often considered impractical as it computes $\log T$ distinct batches and estimators, requiring knowledge of $T$ and, more critically, discards a significant portion of samples at each parameter update. To retain the theoretical efficiency while improving practicality, we adopt the doubly robust (DR) estimation framework, which is known to achieve $\widetilde{O}(\sqrt{dT})$ (Kim et al., 2021, 2023c). Given that $\widetilde{\mathbf{x}}_{a}\in\mathbb{R}^{K}$, we propose an efficient algorithm that employs a DR estimator with ridge regularization and obtains a regret bound of $\widetilde{O}(\sqrt{KT})$ (see Appendix˜C). However, when $K>d$ and $d_{u}=0$, the regret is the regret is higher than that of OLB algorithms. To address this challenge, we propose a novel estimation strategy in the following section that eliminates the dependence on $K$ in the regret bound.
@@ 4.2 Doubly Robust Lasso Estimator
In Eq.˜7, the parameter $\boldsymbol{\mu}_{\star}$ is sparse, with its sparsity depending on the number of nonzero elements required to represent the inaccessible portion of the reward projected onto $\mathrm{R}(\mathbf{X})^{\perp}$, namely $d_{h}$ defined in Eq.˜6. Recall from Eq.˜5 that this projection can be expressed using at most $d_{h}$ basis vectors, implying that $\boldsymbol{\mu}_{\star}^{(u)}$ has at most $d_{h}$ nonzero entries.
@@
$\begin{split}\check{\boldsymbol{\mu}}^{L}_{t}&:=\mathop{\mathrm{argmin}}_{\boldsymbol{\mu}}\sum_{\tau=1}^{t}(y_{a_{\tau},\tau}-\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}\boldsymbol{\mu})^{2}\\ &\qquad+2\widetilde{\sigma}_{\max}\sigma\sqrt{2pt\log\frac{2Kt^{2}}{\delta}}\|\boldsymbol{\mu}\|_{1},\end{split}$ (8)
@@
where $p$ is the coupling probability used to define the multinomial distribution for pseudo-action sampling (as defined in Eq.˜9), and $\delta$ is the confidence parameter of the algorithm. Note that $\widetilde{\sigma}^{2}_{\max}:=\max_{a\in[K]}\mathbf{e}_{a}^{\top}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\mathbf{e}_{a}$, i.e., the largest diagonal entry of the Gram matrix constructed from the augmented feature vectors $\widetilde{\mathbf{x}}_{a}$ over $\mathcal{A}$.
@@
For the estimator in Eq.˜8 to correctly identify the zero entries in $\boldsymbol{\mu}_{\star}$, the compatibility condition needs to be satisfied (van de Geer & Bühlmann, 2009). Although the compatibility condition does not, in general, require a positive minimum eigenvalue for the Gram matrix, in our setting, $\lambda_{\min}(t^{-1}\sum_{s=1}^{t}\widetilde{\mathbf{x}}_{a_{s}}\widetilde{\mathbf{x}}_{a_{s}}^{\top})>0$. Therefore, the compatibility condition is implied without any additional assumption. However, ensuring a sufficiently large minimum eigenvalue typically requires collecting a large number of exploration samples, which in turn increases regret. Achieving this with fewer exploration samples remains a key challenge in the bandit literature, as the minimum eigenvalue affects the convergence rate of the estimator and, consequently, the regret bound (Soare et al., 2014; Kim et al., 2021).
@@
We introduce a doubly robust (DR) estimator that employs the Gram matrix constructed from the augmented feature vectors over the entire action space, $\sum_{s=1}^{t}\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$, rather than only from the chosen actions, $\sum_{s=1}^{t}\widetilde{\mathbf{x}}_{a_{s}}\widetilde{\mathbf{x}}_{a_{s}}^{\top}$. Originating in the missing-data literature (Bang & Robins, 2005), DR estimation yields consistent estimators if either the imputation or the observation probability model is correct (Kim et al., 2021). In the bandit setting, only the reward of the chosen arm is observed for each decision round, leaving the other K1K-1 as missing. Hence, DR estimation imputes these K1K-1 missing rewards and incorporates all the associated feature vectors into the estimation process. Since the observation probabilities are determined by the policy—which is known to the learner—the DR estimator remains robust to errors in the reward estimation. While Kim & Paik (2019) proposed a DR Lasso estimator under IID features satisfying the compatibility condition, we propose a novel DR Lasso estimator that does not rely on such assumptions.
@@
We improve the DR estimation by incorporating resampling and coupling methods. For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote the set of exploration rounds such that for any $\tau\in\mathcal{E}_{t}$, the action $a_{\tau}$ is sampled uniformly over $[K]$. The set $\mathcal{E}_{t}$ is constructed recursively: starting with $\mathcal{E}_{0}=\emptyset$, we define
@@
$\mathcal{E}_{t}=\begin{cases}\mathcal{E}_{t-1}\cup\{t\}&\mathrm{if}\quad|\mathcal{E}_{t}|\leq C_{e}\log\dfrac{2Kt^{2}}{\delta},\\ \mathcal{E}_{t-1}&\mathrm{otherwise}.\end{cases}\vskip-5.69046pt$
@@
Here, $C_{e}$ is defined as $(8K)^{3}\widetilde{\sigma}_{\min}^{-2}\widetilde{\sigma}_{\max}^{2}(1-p)^{-2}$, where $\widetilde{\sigma}_{\min}^{2}:=\lambda_{\min}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})$ denotes the minimum eigenvalue of the Gram matrix constructed from the augmented feature vectors over the entire action space.
@@
At each round $t\notin\mathcal{E}_{t}$, the algorithm selects $a_{t}$ according to an $\epsilon_{t}$-greedy strategy: with probability $\epsilon_{t}=t^{-1/2}$, it selects $\widehat{a}_{t}=\mathop{\mathrm{argmax}}_{a\in[K]}\widetilde{\mathbf{x}}_{a}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}$; with $1-\epsilon_{t}$, it selects an action uniformly at random from $K]\setminus\{\widehat{a}_{t}\}$. Given $a_{t}$, a pseudo-action $\widetilde{a}_{t}$ is sampled from a multinomial distribution:
@@
With the pseudo-actions coupled with the actual actions, we construct unbiased pseudo-rewards for all $a\in[K]$ as: